# Missing Pieces: How Do Designs that Expose Uncertainty Longitudinally Impact Trust in AI Decision Aids? An In Situ Study of Gig Drivers

### v2

@@ Abstract.
Decision aids based on artificial intelligence (AI) induce a wide range of outcomes when they are deployed in uncertain environments. In this paper, we investigate how users’ trust in recommendations from an AI decision aid is impacted over time by designs that expose uncertainty in predicted outcomes. Unlike previous work, we focus on gig driving — a real-world, repeated decision-making context. We report on a longitudinal mixed-methods study ($n=51$) where we measured gig drivers’ trust as they interacted with an AI-based schedule recommendation tool. Our results show that participants’ trust in the tool was shaped by both their first impressions of its accuracy and their longitudinal interactions with it; and that task-aligned framings of uncertainty improved trust by allowing participants to incorporate uncertainty into their decision-making processes. Additionally, we observed that trust depended on their characteristics as drivers, underscoring the need for more in situ studies of AI decision aids.
@@
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; April 23–26, 2025; Athensbooktitle: FAccT ’25: ACM Conference on Fairness, Accountability, and Transparency, April 23–26, 2025, Athensprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing User studiesccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Empirical studies in collaborative and social computingccs: Applied computing Transportationccs: Human-centered computing HCI theory, concepts and models
@@ 1. Introduction
The fairness, accountability, and transparency of algorithms are inextricably intertwined with the extents to which people are willing to interact with them in different contexts. Accordingly, the FAccT community has increasingly focused on studying trust in algorithmic systems, especially those based on artificial intelligence (AI) (Ferrario and Loi, 2022; Jacovi et al., 2021; Kim et al., 2023; Knowles and Richards, 2021; Liao and Sundar, 2022). In this paper, we study how people trust and rely on AI decision aids. AI decision aids function by (1) recommending decisions and (2) predicting how good the outcomes of following those decisions will be. Over recent years, AI decision aids have been adopted at a steadily increasing rate in a number of application domains (Milana and Ashta, 2021; McConvey et al., 2023; Riahi et al., 2021; Skjuve et al., 2023; Wang et al., 2021; Završnik, 2019). However, because AI decision aids are used in the context of people and other sociotechnical systems, interactions with these entities can lead to considerable uncertainty in outcomes (Russell et al., 2017; Macrae, 2021). When this uncertainty impacts the predictability of AI decision aids, users’ trust in the decision aid can be eroded (Jacovi et al., 2021; Solberg et al., 2022).
Prior literature on trust in AI decision aids under uncertainty can be organised into two complementary lines of work. One line of work has studied specific factors that influence trust in AI decision aids, using laboratory experiments in simulated, single-shot, and low-stakes scenarios that require limited domain expertise (Bansal et al., 2021; Cai et al., 2019; Jacobs et al., 2021; Kocielnik et al., 2019; Zhang et al., 2020). However, context is an important factor for trust in AI (Kim et al., 2023; Solberg et al., 2022). Therefore, the contrived nature of these experiments limits their generalisability to the real-world use contexts of AI decision aids. Another line of work has studied trust in real-world AI decision aids (Beede et al., 2020; Kim et al., 2023; Wang et al., 2021), using observational, qualitative studies to assess how existing users interact with these decision aids. Despite their ecological validity, these studies do not quantitatively assess design factors and thus provide limited insight into how new, trustworthy AI decision aids can be designed.
@@ 2.1. Trust in AI
A lack of uniformity exists in the human-AI interaction literature on how to define and evaluate trust in AI systems (Ueno et al., 2022). We follow Mayer et al. (1995) in operationalising the constructs of trust and reliance: in the context of AI decision aids, they hypothesise that trust is “the willingness of a person to be vulnerable to the actions of an [AI decision aid], based on the expectation that it will perform a [decision-making task] important to the trustor”, and that reliance is the external, behavioural expression of that internal attitude (Solberg et al., 2022). In accordance with this definition, one point of broad consensus in the literature is that the design of AI systems impacts expectations of their performance in decision-making, and thus the trust of their users (Burgess et al., 2023; Danner et al., 2020; Jacobs et al., 2021; Kocielnik et al., 2019; Kunkel et al., 2019; Scharowski et al., 2023; Yu et al., 2019).
However, the majority of this work has been based on controlled laboratory experiments. Compared to real-world use contexts of AI decision aids, such experiments have two main limitations: (1) They are single-shot, involving only a single session of AI use with no temporal separation between decision points (Ueno et al., 2022). Several studies have compared two trust measurements (Mou and Cohen, 2017; Mou et al., 2017; Kunkel et al., 2019), but these studies still provide a limited perspective on the evolution of trust dynamics. (2) They are low-stakes, involving contrived or hypothetical decision-making scenarios in which an element of risk and thus vulnerability is largely absent (Kim et al., 2023). Some work has studied AI decision aids for decision-making domains entailing high stakes in the real world, such as the medical (Burgess et al., 2023; Cai et al., 2019; Beede et al., 2020; Jacobs et al., 2021; Wang et al., 2021; Sivaraman et al., 2023; Wysocki et al., 2023; Yang et al., 2023) and financial (Prabhudesai et al., 2023; Schoeffer et al., 2022; Scharowski et al., 2023) contexts. However, for practical and ethical reasons, participants in these studies cannot receive feedback from the real world for their decisions. A minority of work has evaluated trust in AI decision aids within their real-world contexts (Beede et al., 2020; Kim et al., 2023; Wang et al., 2021; Widder et al., 2021), but these have been limited to observational studies that did not compare multiple designs.
@@ 2.2. Gig Work and Gig Driving
Gig work offers workers the flexibility and autonomy to choose when and where they want to work (Lee et al., 2015). However, this autonomy is hampered by the opacity of platforms’ assignment, pricing, and evaluation mechanisms (Griesbach et al., 2019; Lee et al., 2015; Singh et al., 2023; Tan et al., 2021; Wood et al., 2019). In particular, gig driving platforms impose information asymmetry (Maffie, 2023; Rosenblat and Stark, 2016) by dynamically varying their compensation mechanisms (Lee et al., 2015; Zhang et al., 2022a). For drivers, this lack of visibility leads to difficulties in planning (Griesbach et al., 2019) and inequities in revenue (Miao et al., 2023). This makes gig driving an exemplary context under which the impact of uncertainty, repeated interactions, and moderate financial stakes on AI decision aids can be assessed.
Within the gig driving setting, the most relevant prior work to our paper consists of studies that have designed predictive and prescriptive tools for gig drivers (Battifarano and Qian, 2019; Khan et al., 2022; Zhang et al., 2022a, 2023, 2024). Among these, Battifarano and Qian (2019); Khan et al. (2022) focused on evaluating the predictive accuracy of their systems and did not include a user design component; Zhang et al. (2022a, 2023) were formative studies that designed tools in collaboration with drivers but did not deploy them to evaluate users’ trust and reliance; and Zhang et al. (2024) tested the effects of different AI decision aid designs that displayed uncertainty, albeit using simulated taxi trip data and focusing on the decisions — not attitudes — of participants.
@@ 3. AI-Based Schedule Recommendation for Gig Driving
In this section, we describe an AI-based schedule recommendation tool for gig drivers. Like other AI decision aids (Solberg et al., 2022), this tool (1) recommends a set of decisions (i.e. a schedule) to achieve a set objective (see Section 3.1), and (2) predicts the outcomes (i.e. estimated earnings) of following those decisions. We use this tool as an exemplary AI decision aid to study the longitudinal relationship between the framing of outcome uncertainty and trust. The design of this tool was conceptualised and refined through a formative study, which we describe in Appendix B.
@@ 3.1. Decision Aid Design
Our tool consists of two modules. First, an estimation module prospectively predicts $e_{ij}$, the earnings that drivers can expect during a specific hour $j$ on a specific weekday $i$. These could be computed by a machine learning model or averaged from historical data. Second, for each driver, a scheduling module uses the estimated earnings and the driver’s constraints as inputs to produce an optimal set of working times. To do so, it solves a constrained optimisation problem to set variables $x_{ij}$ to 1 or 0, denoting whether the driver is recommended to work during each time slot $(i,j)$.
@@
  • Some drivers who value earnings to a greater extent set personal minimum targets for their hourly or daily earnings instead of restricting their driving hours. For these drivers, we minimise the objective function: the total hours of recommended timeslots throughout the week, i.e. $\sum_{i,j}x_{ij}$.
@@ 3.3. Interface Conditions
The schedule page uses the outputs of the estimation module to predict how much a driver following the recommended schedule would make per hour or per week. However, uncertainty inherently exists in these predictions, as they are based on historical data and their realisation is contingent upon which gigs are offered to and accepted by drivers. To address Research Question 2, we varied the design of the schedule page between four conditions (Figure 1):
@@
    Base condition. Users were only shown their mean weekly estimated earnings, and their mean estimated earnings for each hour during the week.
@@
    Ranged estimates. To assess the effect of exposing uncertainty through range-based estimates (similar to (Prabhudesai et al., 2023)), users were shown mean, pessimistic, and optimistic estimates for hourly and weekly earnings.
@@
    Ranged and hedged estimates. To assess the effect of exposing uncertainty through lexical hedging (similar to (Kim et al., 2024; Zhang et al., 2022b)), the textual description of the estimates was changed from (R). Instead of “Based on historical data, it is estimated that you will earn”, (RH) states “On average, based on historical data, a driver following this schedule will earn”.
@@ 4.1. Participants and Data Sources
Accordingly, we generated estimates using gig data from August 2023 in each of these four cities. For each city, our data included approximately 100 000–300 000 gig records distributed evenly across times and weekdays. Hourly earnings were estimated by the mean of what drivers historically earned in this slot, filtered to the participant’s city and platforms. We used this static estimator to focus on the effects of exposing uncertainty in the estimates. Thus, we did not emphasise to participants that the schedule page was generated using AI or optimisation, and did not provide any additional information about the data used to generate the estimates.
@@ 4.2. User Study Activities
Participants received a link to the study website from an email-distributed recruitment message. After the consent form, they completed the first of four surveys, the Intake Survey (Section A.1). This 12-question survey asked about their needs and motivations as gig drivers, along with demographics.
Next, participants were directed to interact with the tool, which we displayed in an iframe to mitigate response bias (Dell et al., 2012; Ko et al., 2015). They entered their constraints on the constraint page, and received the tool’s recommended schedule for the entire week on the schedule page. Participants were randomly and uniformly assigned to one of the four conditions for the schedule page (Section 3.3), such that each condition had an approximately equal number of participants.
@@
At the end of each participant’s indicated availability, a second scheduled email sent them a link to the End-of-Day Survey (Section A.3). This was an 8-question survey that measured their trust in the tool for that day, and their intention to rely on the tool for the next day (Section 5.1). If the participant intended to continue relying on the tool, they were then presented with a daily variant of the schedule page. Here, they could review the recommended schedule for the following day, and revise their constraints for the day as desired. Updated schedules were generated by fixing the recommended time slots for previous days using equality constraints and then re-solving the optimisation problem. However, if the participant intended to pause their interaction for one day, an email was sent on the next day, which prompted them to either review the next day’s schedule or to pause for an additional day.
@@
On the final day, we removed the last question of the End-of-Day Survey and added the Post-Survey (Section A.4). This was a 10-question survey that retrospectively measured participants’ trust and distrust in the tool over the course of the entire user study (Section 5.1). After completing the Post-Survey, participants were sent a final email that invited them to participate in an optional Exit Interview (Section 4.3).
@@ 5. Quantitative Analysis
Out of these 51 participants, 44 completed at least one day of interaction with the tool, and 34 completed all 7 days of interaction. Starting from Day 0, Day 7 was reached by 6 (46%) of the baseline Condition (B) participants; 10 (71%) of the daily estimate Condition (D) participants; 7 (58%) of the ranged estimate Condition (R) participants; and 11 (92%) of the ranged and hedged Condition (RH) participants. We show full retention statistics in Figure 3.

### v4

@@ Abstract.
Decision aids based on artificial intelligence (AI) induce a wide range of outcomes when they are deployed in uncertain environments. In this paper, we investigate how users’ trust in recommendations from an AI decision aid is impacted over time by designs that expose uncertainty in predicted outcomes. Unlike previous work, we focus on gig driving — a real-world, repeated decision-making context. We report on a longitudinal mixed-methods study ($n=51$) where we measured gig drivers’ trust as they interacted with an AI-based schedule recommendation tool. Our results show that participants’ trust in the tool was shaped by both their first impressions of its accuracy and their longitudinal interactions with it, and that task-aligned framings of uncertainty improved trust by allowing participants to incorporate uncertainty into their decision-making processes. Additionally, we observed that trust depended on their characteristics as drivers, underscoring the need for more in situ studies of AI decision aids.
@@
journalyear: 2025copyright: ccconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732050isbn: 979-8-4007-1482-5/2025/06ccs: Human-centered computing User studiesccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Empirical studies in collaborative and social computingccs: Applied computing Transportationccs: Human-centered computing HCI theory, concepts and models
@@ 1. Introduction
The fairness, accountability, and transparency of algorithms are inextricably intertwined with the extents to which people are willing to interact with them. Accordingly, the FAccT community has increasingly focused on studying trust in algorithmic systems, especially those based on artificial intelligence (AI) (Ferrario and Loi, 2022; Jacovi et al., 2021; Kim et al., 2023; Knowles and Richards, 2021; Liao and Sundar, 2022). In this paper, we study how people trust and rely on AI decision aids. AI decision aids function by (1) recommending decisions and (2) predicting how good the outcomes of following those decisions will be. Over recent years, AI decision aids have been adopted at an increasing rate in a number of application domains (Milana and Ashta, 2021; McConvey et al., 2023; Riahi et al., 2021; Skjuve et al., 2023; Wang et al., 2021; Završnik, 2019). However, because AI decision aids are used in the context of people and other sociotechnical systems, considerable uncertainty in decision outcomes can exist (Russell et al., 2017; Macrae, 2021). When this uncertainty impacts the predictability of AI decision aids, users’ trust in the decision aid can be eroded (Jacovi et al., 2021; Solberg et al., 2022).
Prior literature on trust in AI decision aids under uncertainty can be organised into two complementary lines of work. One line of work has studied specific factors that influence trust in AI decision aids, using laboratory experiments in simulated, single-shot, and low-stakes scenarios that require limited domain expertise (Bansal et al., 2021; Cai et al., 2019; Jacobs et al., 2021; Kocielnik et al., 2019; Zhang et al., 2020). However, context is an important factor for trust in AI (Kim et al., 2023; Solberg et al., 2022). Therefore, the contrived nature of these experiments limits their generalisability to the real-world use contexts of AI decision aids. Another line of work has studied trust in real-world AI decision aids (Beede et al., 2020; Kim et al., 2023; Wang et al., 2021), using observational, qualitative studies to assess how existing users interact with these decision aids. These studies are not quantitative assessments of design factors and provide limited insight into new, trustworthy AI decision aids can be designed.
@@ 2.1. Trust in AI
A lack of uniformity exists in the human-AI interaction literature on how to define and evaluate trust in AI systems (Ueno et al., 2022). We follow Mayer et al. (1995) in operationalising the constructs of trust and reliance: in the context of AI decision aids, they hypothesise that trust is “the willingness of a person to be vulnerable to the actions of an [AI decision aid], based on the expectation that it will perform a [decision-making task] important to the trustor”, and that reliance is the external, behavioural expression of that internal attitude (Solberg et al., 2022). One point of broad consensus in the literature is that the design of AI systems impacts expectations of their performance in decision-making, and thus the trust of their users (Burgess et al., 2023; Danner et al., 2020; Jacobs et al., 2021; Kocielnik et al., 2019; Kunkel et al., 2019; Scharowski et al., 2023; Yu et al., 2019).
However, the majority of this work has been based on controlled laboratory experiments. Compared to real-world use contexts of AI decision aids, such experiments have two main limitations: (1) They are single-shot, involving only a single session of AI use with no temporal separation between decision points (Ueno et al., 2022). Several studies have compared two trust measurements (Mou and Cohen, 2017; Mou et al., 2017; Kunkel et al., 2019), but these studies still provide a limited perspective on the evolution of trust dynamics. (2) They are low-stakes, involving contrived or hypothetical decision-making scenarios in which an element of risk and thus vulnerability is largely absent (Kim et al., 2023). Some work has studied AI decision aids for decision-making domains that entail high stakes in the real world, such as the medical (Burgess et al., 2023; Cai et al., 2019; Beede et al., 2020; Jacobs et al., 2021; Wang et al., 2021; Sivaraman et al., 2023; Wysocki et al., 2023; Yang et al., 2023) and financial (Prabhudesai et al., 2023; Schoeffer et al., 2022; Scharowski et al., 2023) contexts. However, for practical and ethical reasons, participants in these studies cannot receive feedback from the real world for their decisions. A minority of work has evaluated trust in AI decision aids within their real-world contexts (Beede et al., 2020; Kim et al., 2023; Wang et al., 2021; Widder et al., 2021), but these have been limited to observational studies that did not compare multiple designs.
@@ 2.2. Gig Work and Gig Driving
Gig work offers workers the flexibility and autonomy to choose when and where they want to work (Lee et al., 2015). However, this autonomy is hampered by the opacity of platforms’ assignment, pricing, and evaluation mechanisms (Griesbach et al., 2019; Lee et al., 2015; Singh et al., 2023; Tan et al., 2021; Wood et al., 2019). In particular, gig driving platforms impose information asymmetry (Maffie, 2023; Rosenblat and Stark, 2016) by dynamically varying their compensation mechanisms (Lee et al., 2015; Zhang et al., 2022a). For drivers, this lack of visibility leads to difficulties in planning (Griesbach et al., 2019) and inequities in revenue (Miao et al., 2023). This makes gig driving an exemplary context to assess the impact of uncertainty, repeated interactions, and moderate financial stakes on AI decision aids.
Within the gig driving setting, the most relevant prior work to our paper consists of studies that have designed predictive and prescriptive tools for gig drivers (Battifarano and Qian, 2019; Khan et al., 2022; Zhang et al., 2022a, 2023, 2024). Among these, Battifarano and Qian (2019) and Khan et al. (2022) focused on evaluating the predictive accuracy of their systems and did not include a user design component; Zhang et al. (2022a, 2023) conducted formative studies that designed tools in collaboration with drivers but did not deploy them to evaluate users’ trust and reliance; and Zhang et al. (2024) tested the effects of different AI decision aid designs that displayed uncertainty, albeit using simulated taxi trip data and focusing on the decisions — not attitudes — of participants.
@@ 3. AI-Based Schedule Recommendation for Gig Driving
In this section, we describe an AI-based schedule recommendation tool for gig drivers. Like other AI decision aids (Solberg et al., 2022), this tool (1) recommends a set of decisions (i.e. a schedule) to achieve a set objective (see Section 3.1), and (2) predicts the outcomes (i.e. estimated earnings) of following those decisions. We use this tool as an exemplary AI decision aid to study the longitudinal relationship between the framing of uncertainty in outcomes and trust. The design of this tool was conceptualised and refined through a formative pilot study, which we describe in Appendix B.
@@ 3.1. Decision Aid Design
Our tool consists of two modules. First, an estimation module prospectively predicts $e_{ij}$, the earnings that drivers can expect during a specific hour $j$ on a specific weekday $i$. These could be computed by a machine learning model or averaged from historical data. Second, for each driver, a scheduling module uses the estimated earnings and the driver’s constraints as inputs to produce an optimal set of working times. To do so, it solves a constrained optimisation problem to set variables $x_{ij}$ to 1 or 0, denoting whether the driver is recommended to work in time slot $(i,j)$.
@@
  • Some drivers who value earnings to a greater extent set minimum targets for their hourly or daily earnings instead of restricting their driving hours. For these drivers, we minimise the objective function: the total hours of recommended time slots throughout the week, i.e. $\sum_{i,j}x_{ij}$.
@@ 3.3. Interface Conditions
The schedule page uses the outputs of the estimation module to predict how much a driver following the recommended schedule would make per hour and per week. However, uncertainty inherently exists in these predictions, as they are based on historical data, and their realisation is contingent upon which gigs are offered to and accepted by drivers. To address Research Question 2, we varied the design of the schedule page between four conditions (Figure 1):
@@
    Base condition. Users were only shown their mean estimated earnings for the week and for each hour in the week.
@@
    Ranged estimates. To assess the effect of exposing uncertainty through range-based estimates (similar to Prabhudesai et al. (2023)), users were shown mean, pessimistic, and optimistic estimates for hourly and weekly earnings.
@@
    Ranged and hedged estimates. To assess the effect of exposing uncertainty through lexical hedging (similar to Kim et al. (2024) and Zhang et al. (2022b)), the textual description of the estimates was changed from (R). Instead of “Based on historical data, it is estimated that you will earn”, (RH) states “On average, based on historical data, a driver following this schedule will earn”.
@@ 4. Longitudinal User Study Design
[…]
Screenshots of a box titled ”Shift Recommendation Tool” on the schedule page. Condition (B) is shown at the top. An estimated weekly earning of \$1412 is highlighted with box 1. A recommended schedule shows average earnings for each weekday and each hour. Two cells reading ”Average: \$15” and ”Average: \$12” are highlighted with box 2. Before the weekly estimate of \$1412, box 3 highlights text reading ”Based on historical data, it is estimated that you will earn a weekly average of”. The subsequent three subfigures show variants of this page. For box 1 (weekly earnings), Condition (D) displays a table showing ”An average of…” with \$210 for Monday, \$194 for Tuesday, etc.; Conditions (R) and (RH), a table showing ”On an average week: \$1412”, ”On a bad week: \$391”, and ”On a good week: \$3032”. For box 2 (hourly earnings), Conditions (R) and (RH) display cells reading ”Average: \$15, Low: \$4, High: \$34” and ”Average: \$12, Low: \$4, High: \$28”. For box 3 (textual description), Condition (RH) displays ”In general, every driver may encounter different conditions while they are driving, which leads to differences in earnings from individual to individual and from one day to the next. On average, based on historical data, a driver following this schedule will earn”.
@@
[…]
Flow chart showing bulleted lists of study activities. On Day 0, the Intake Survey is followed by use of the tool to generate the weekly schedule, along with the Pre-Survey as a first measurement of trust. On Days 1-6, the participant receives a daily schedule, drives as usual, and completes the End-of-Day Survey. This repeats until Day 7, which is similar except the End-of-Day Survey is followed by the Post-Survey. This is followed by the optional exit interview.
@@ 4.1. Participants and Data Sources
Accordingly, we generated estimates using gig data from August 2023 in each of these four cities. For each city, our data included approximately 100 000–300 000 gig records distributed evenly across times and weekdays. Hourly earnings were estimated by the mean of what drivers historically earned in this slot, filtered to the participant’s city and platforms. We used this static estimator to focus on the effects of exposing uncertainty in the estimates. Accordingly, we did not emphasise to participants that the schedule page was generated using AI or optimisation, and did not provide any additional information about the data used to generate the estimates.
@@ 4.2. User Study Activities
Participants received a link to the study website from a recruitment message distributed by email. After the consent form, they completed the first of four surveys, the Intake Survey (Section A.1). This 12-question survey asked about their needs and motivations as gig drivers, along with demographics.
Next, participants were directed to interact with the tool, which we displayed in an iframe to mitigate response bias (Dell et al., 2012; Ko et al., 2015). They entered their constraints on the constraint page, and received the tool’s recommended schedule for the entire week on the schedule page. Participants were assigned to one of the four conditions for the schedule page (Section 3.3) uniformly at random, such that each condition had an approximately equal number of participants.
@@
At the end of each participant’s indicated for the day availability, a second scheduled email sent them a link to the End-of-Day Survey (Section A.3). This was an 8-question survey that measured their trust in the tool for that day, and their intention to rely on the tool for the next day (Section 5.1). If the participant intended to continue relying on the tool, they were then presented with a daily variant of the schedule page. Here, they could review the recommended schedule for the following day, and revise their constraints for the day as desired. Updated schedules were generated by fixing the recommended time slots for previous days using equality constraints and then re-solving the optimisation problem. However, if the participant intended to pause their interaction for one day, an email was sent on the next day, which prompted them to either review the next day’s schedule or to pause for an additional day.
@@
On the final day, we removed the last question measuring reliance from the End-of-Day Survey, and added the Post-Survey (Section A.4). This was a 10-question survey that retrospectively measured participants’ trust and distrust in the tool over the entire user study (Section 5.1). After completing the Post-Survey, participants were sent a final email that invited them to participate in an optional Exit Interview (Section 4.3).
@@ 5. Quantitative Analysis
Out of these 51 participants, 44 completed at least one day of interaction with the tool, and 34 completed all 7 days of interaction. Starting from Day 0, Day 7 was reached by 6 (46%) of the base Condition (B) participants; 10 (71%) of the daily estimate Condition (D) participants; 7 (58%) of the ranged estimate Condition (R) participants; and 11 (92%) of the ranged and hedged Condition (RH) participants. We show full retention statistics in Figure 3.
@@
[…]
Bar graph showing the number of participants from Day 0 to Day 7, followed by the Interview. On Day 0, the four conditions respectively had 13, 14, 12, and 12 participants. Over the following week, retention for Conditions (D) and (RH) remained high, but retention dropped substantially for Conditions (B) and (R). By Day 7, the four conditions respectively had 6, 10, 7, and 11 participants. The Interview had 0, 3, 1, and 3 participants for the four conditions.